OpenStudy (anonymous):

F(x) = The integral from 15 to x of sqrt(t^2 + 64) dt. Evaluate at F(15) and F'(15)

OpenStudy (anonymous):

\[\int\limits_{15}^{x} \sqrt{t ^{2}+64} dt\]

OpenStudy (anonymous):

\[F(15)=\int _{15}^{15}\sqrt{t^2+64}dt=0\] because the "length of the path" is zero

OpenStudy (anonymous):

it is always the case that \[\int_a^af(t)dt=0\]

OpenStudy (anonymous):

Ok, second part: I'm not that great with integrals and I'm unsure how to integrate the function. Using u-sub for t^2 + 64 would get du of 2t. But since there is no t outside the function, I'm not quite sure what to do...

OpenStudy (anonymous):

you are not asked to integrate the function don't worry about finding a nice closed form for the integral you are only asked for \(F(15)\) and \(F'(15)\)

OpenStudy (anonymous):

it is a common misunderstanding that given some function \(f(tx)=\text{whatever}\) then somehow one can find some other function in a nice closed form \(F(x)\) with \(F'(x)=f(x)\)

OpenStudy (anonymous):

it is almost never the case

OpenStudy (anonymous):

Ok. So I get that F(15) = 0. You can't have a length or area under a distance of nothing. However, what does it mean when it asks F'(15) ? What should I be looking to do at that point?

OpenStudy (anonymous):

the derivative of the integral is the integrand

OpenStudy (anonymous):

So it should be \[\sqrt{t^{2}+64}\], yeah? At least, I already tried that but the program says it is incorrect

OpenStudy (anonymous):

well actually it should be \[\sqrt{x^2+64}\] since \[F(x)=\int _{15}^x\sqrt{t^2+64}dt\] is a function of \(x\) not \(t\) so \(F'(x)=\sqrt{x^2+64}\)

OpenStudy (anonymous):

But it's F'(15)

OpenStudy (anonymous):

you were asked not for \(F'(x)\) but rather \(F'(15)\) right?

OpenStudy (anonymous):

Yes. I get that x replaces t when it is asked as a function of x. Just not quite sure when its 15

OpenStudy (anonymous):

i think you got confused maybe because you were thinking of integrals suppose you were in a basic pre - calc course and were given the problem as follows : if \(f(x)=\sqrt{x^2+64}\) what is \(f(15)\)? i bet you could solve in seconds

OpenStudy (anonymous):

Bleh. 17.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I was. One last one. so if it says F"(15). Take the derivative of f(x) and evaluate at 15?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

lol, thanks. If you have a moment, may I ask another question?

OpenStudy (anonymous):

\[\frac{x}{\sqrt{x^2+64}}\] is what i get

OpenStudy (anonymous):

i mean as the second derivative at 15 i guess you get \(\frac{15}{17}\)

OpenStudy (anonymous):

I got the same

OpenStudy (anonymous):

yeah go ahead and ask

OpenStudy (anonymous):

I'm working on finding a total distance traveled for rectilinear motion. \[v(t) = t ^{3} - 3t ^{2} + 2t; 0\le t \le 5\]

OpenStudy (anonymous):

I get \[s(t) = t ^{4}/4 -t ^{3} + t ^{2}\]

OpenStudy (anonymous):

And I got displacement as equal to 225/4 which worked out. But I'm unsure how to find total distance traveled

OpenStudy (anonymous):

be careful here because my guess is you have to break up the integral in to parts where it is going forward and going backwards

OpenStudy (anonymous):

For the equation I found that t = 2 is where it goes from negative to positive. However, I don't know what to do with that information

OpenStudy (anonymous):

integrate from 0 to 2, then from 2 to 5 first one is positive, second one is negative, then add the absolute values i did not check that 2 was right, but i believe you

OpenStudy (anonymous):

But if 2 is zero and 0 is zero, then the integral from 0 to 2 is just zero?

OpenStudy (anonymous):

Then 2 to 5 is the same as 0 to 5 because its still 5 - 0

OpenStudy (anonymous):

hold on one second, let me check this

OpenStudy (anonymous):

I think this is where I'm getting pretty confused

OpenStudy (anonymous):

ok first off, it is \[f(t)=t^3-3t^2+2t\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

which by some happy coincidence is \(f(t)=t(t-1)(t-2)\)

OpenStudy (anonymous):

so in fact unfortunately for you it changes sign at \(1\) and at \(2\)

OpenStudy (anonymous):

Evaluate at the base function not at the integral

OpenStudy (anonymous):

on the interval \([0,5]\) it is positive, then negative, then positive

OpenStudy (anonymous):

First problem found

OpenStudy (anonymous):

yeah hold on it is the "speed" except it is actually "velocity" which can be negative you are in your car driving forward for one hour, then backwards for one hour, then forwards for 3 hours your "displacement" is just how far you are from home, but the total distance travelled is the forwards distance plus the backwards distance plus the forwards distance

OpenStudy (anonymous):

So I get that from 0 to 1, it is 1/4, 1 to 2 it is 1/4 again(taking abs value) and then my original value of 225/4 from 2 to 5. Does this seem about right?

OpenStudy (anonymous):

well i didn't to it, but that is certainly the right idea

OpenStudy (anonymous):

Yeah, I just didn't see how I was messing up because I knew I needed separate values to add together showing total forward/backward distance

OpenStudy (anonymous):

Perfect, thanks again!

OpenStudy (anonymous):

yw